Question 1

Calculate the pH of a 1.0 × 10−6M solution of HNO3

Question 2

Calculate the pH of a 1.0 × 10−1M solution of NaOH
Answer #1

1)

[H+] = [HNO3] = 1.0*10^-6 M

use:

pH = -log [H+]

= -log (1.0*10^-6)

= 6.00

Answer: 6.00

2)

[OH-] = [NaOH] = 1.0*10^-1 M

use:

pOH = -log [OH-]

= -log (1.0*10^-1)

= 1.00

use:

PH = 14.00 - pOH

= 14.00 - 1.00

= 13.00

Answer: 13.00